A company has beginning inventory of 32 units at a cost of $12.00 each on October 1. On October 5, it purchases 22 units at $13.00 per unit. On October 12 it purchases 32 units at $14.00 per unit. On October 15, it sells 66 units. Using the FIFO periodic inventory method, what is the value of the inventory at October 15 after the sale?
Answer #1
Unit Unit Cost Total cost
Beginning inventory 32 12 384
Oct 5 22 13 286
Oct 12 32 14 448
Total 86

Ending inventory Unit = 86-66 = 20 Units

Value ending inventory = 20*14 = $280